A short quiz has two true-false questions and one multiple-choice question with four possible answers. A student guesses at each question. Assuming the choices are all equally likely and the questions are independent of each other, the following is the probability distribution of the number of answers guessed correctly.

What is the Probability of getting less than all three right?

X            0            1            2            3

P(X)       .1875    .4375    .3125    .0625

Group of answer choices

a) 1.0000

b) .9375

c) .6250

d) .3125
Answer #1

P(X < 3) = P(X = 0) + P(X = 1) + P(X = 2)
= 1 - P(X = 3)
= 1 - 0.0625
= 0.9375

option B